how many grams of KCl would have to be diisolved in 60 gms H2O to give 20% by weight of solution?

Grade:12

2 Answers

Naveen Kumar
9 years ago
total mass of the water=60g
let weight of KCl required=xg
so weight %=20%
i.e. {x/(x+60)}=20%=20/100
x=0.2*x +....60*0.2
0.8*x=12
x=12/0.8=15g
.….!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!.................................................................
Naveen Kumar
9 years ago
let amount=xg
so x/(x+60)=20%=0.2
x=15g.....
so the amount of KCL needed=15g
..................................
......................................
…...................................................
/.........///////////////////////////////////////////